ME Text Direction Layout Reset section Cut -, Shape Fill· Fi Align Ten Convert to SmartAr Shape Outline, an Arrange Quick Shape Effectss Copy . aste New { } ☆-)Arrange Quick Stides Font Paragraph Existing Process S stations-3 manual, 2 automatic 1 Supervisor 50 units/hr @104,000 units/yr Proposed Process 5 stations 2 manual,3 automatic 1 Supervisor 50 units/hr 104,000 unts/ Manual 2 oper Auto Auto Auto Auto Auto 1 oper 2 oper St 2 1.25/unit$0.75 25/yr 1 oper Sts 0 2 oper St3 St4 Costs St 1 RawM $1.20/unit $2.75/unit Indirect M $0.25 Spare parts o Tools/Fix $100/y Utilities 0 50 $5,000/vS0.10 $350/yr s0.35/unit 3,500/yr 500/r 1.00/unit Proposed Process A capital investment of 117,000 would purchase the piece of automation which does t process as the current operators and eliminate the need for 2 operators at station 2 the same is this a good project? Direct Labor $15.50/hr Indirect Labor $18.00/hr Supervisor $50,000/yr 1 Maintenance Oper $45,000/ 1 Fork Lift oper $40,000/yr 2
Answer #1

The approach for solving this question would be straight forward - we will compute the cost in both the options and compare with each other. whichever comes lesser will be accepted.

Total cost in Option 1 :

Total Cost A. Operators B. Units per year C. Raw Material/unit D. Indirect Mat/unit E. Utilities/unit F. Total per unit cost

W. note 1 - Calculation of total hours : It is given that 50 units are made in 1 hour and total number of units in the year are 104,000. which simply means that total number of hrs taken are - 104,000/50 units per hr = 2,080 hrs.

Thus, direct labour cost above is 15.50x2,080 = $32,240. On same lines, indirect labour has also been calculated.

so the total cost in option 1 is 1,061,755.

Option 2 : All the cost will be same except the operators cost. The 2 operators will be eliminated and we can save the cost of 40,000+45,000 = 85,000 per year. Also, we will have to make an extra capital investment of 117,000, the useful life of this asset is not given in the question.

Proposed Plan Cost at current option Less: Elimination of 2 operators Net Cost 1,061,755 (85,000) 976,755

So the net cost after eliminating the cost of 2 operators is 976,755. Now we have also made an extra investment of 117,000. The useful life is not given in the question but as it is a capital investment (asset) the useful life must be more than one year say 2 years (assumption).

There, the per year cost of investment written off will be 117,000/2 = 58,500.

Thus, the net cost would be 976,755+58,500 = 1,035,255. which is lower than cost in option 1 thus it is acceptable.

shout if there is any doubt.